\begin{thebibliography}{40}
\expandafter\ifx\csname natexlab\endcsname\relax\def\natexlab#1{#1}\fi

\bibitem[{Angelova et~al.(2022)Angelova, Avramidis, and
  M{\"o}ller}]{angelova-etal-2022-using}
Galina Angelova, Eleftherios Avramidis, and Sebastian M{\"o}ller. 2022.
\newblock \href {https://doi.org/10.18653/v1/2022.acl-srw.21} {Using neural
  machine translation methods for sign language translation}.
\newblock In \emph{Proceedings of the 60th Annual Meeting of the Association
  for Computational Linguistics: Student Research Workshop}, pages 273--284,
  Dublin, Ireland. Association for Computational Linguistics.

\bibitem[{Bawden et~al.(2018)Bawden, Sennrich, Birch, and
  Haddow}]{bawden-etal-2018-evaluating}
Rachel Bawden, Rico Sennrich, Alexandra Birch, and Barry Haddow. 2018.
\newblock \href {https://doi.org/10.18653/v1/N18-1118} {Evaluating discourse
  phenomena in neural machine translation}.
\newblock In \emph{Proceedings of the 2018 Conference of the North {A}merican
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, Volume 1 (Long Papers)}, pages 1304--1313, New Orleans,
  Louisiana. Association for Computational Linguistics.

\bibitem[{Camg{\"o}z et~al.(2018)Camg{\"o}z, Hadfield, Koller, Ney, and
  Bowden}]{8578910}
Necati~Cihan Camg{\"o}z, Simon Hadfield, Oscar Koller, Hermann Ney, and Richard
  Bowden. 2018.
\newblock \href {https://doi.org/10.1109/CVPR.2018.00812} {Neural sign language
  translation}.
\newblock In \emph{2018 IEEE/CVF Conference on Computer Vision and Pattern
  Recognition}, pages 7784--7793.

\bibitem[{Camg{\"o}z et~al.(2020{\natexlab{a}})Camg{\"o}z, Koller, Hadfield,
  and Bowden}]{camgoz2020b}
Necati~Cihan Camg{\"o}z, Oscar Koller, Simon Hadfield, and Richard Bowden.
  2020{\natexlab{a}}.
\newblock \href {https://doi.org/10.1007/978-3-030-66823-5_18} {Multi-channel
  transformers for multi-articulatory sign language translation}.
\newblock In \emph{Computer Vision – ECCV 2020 Workshops: Glasgow, UK, August
  23–28, 2020, Proceedings, Part IV}, page 301–319, Berlin, Heidelberg.
  Springer-Verlag.

\bibitem[{Camg{\"o}z et~al.(2020{\natexlab{b}})Camg{\"o}z, Koller, Hadfield,
  and Bowden}]{camgoz2020sign}
Necati~Cihan Camg{\"o}z, Oscar Koller, Simon Hadfield, and Richard Bowden.
  2020{\natexlab{b}}.
\newblock Sign language transformers: Joint end-to-end sign language
  recognition and translation.
\newblock In \emph{IEEE Conference on Computer Vision and Pattern Recognition
  (CVPR)}.

\bibitem[{De~Coster et~al.(2022)De~Coster, Shterionov, Van~Herreweghe, and
  Dambre}]{de2022machine}
Mathieu De~Coster, Dimitar Shterionov, Mieke Van~Herreweghe, and Joni Dambre.
  2022.
\newblock Machine translation from signed to spoken languages: State of the art
  and challenges.
\newblock \emph{arXiv preprint arXiv:2202.03086}.

\bibitem[{Denkowski and Neubig(2017)}]{denkowski-neubig-2017-stronger}
Michael Denkowski and Graham Neubig. 2017.
\newblock \href {https://doi.org/10.18653/v1/W17-3203} {Stronger baselines for
  trustable results in neural machine translation}.
\newblock In \emph{Proceedings of the First Workshop on Neural Machine
  Translation}, pages 18--27, Vancouver. Association for Computational
  Linguistics.

\bibitem[{Ding et~al.(2019)Ding, Renduchintala, and Duh}]{ding-etal-2019-call}
Shuoyang Ding, Adithya Renduchintala, and Kevin Duh. 2019.
\newblock \href {https://aclanthology.org/W19-6620} {A call for prudent choice
  of subword merge operations in neural machine translation}.
\newblock In \emph{Proceedings of Machine Translation Summit XVII: Research
  Track}, pages 204--213, Dublin, Ireland. European Association for Machine
  Translation.

\bibitem[{Egea~G{\'o}mez et~al.(2021)Egea~G{\'o}mez, McGill, and
  Saggion}]{egea-gomez-etal-2021-syntax}
Santiago Egea~G{\'o}mez, Euan McGill, and Horacio Saggion. 2021.
\newblock \href {https://aclanthology.org/2021.bucc-1.4} {Syntax-aware
  transformers for neural machine translation: The case of text to sign gloss
  translation}.
\newblock In \emph{Proceedings of the 14th Workshop on Building and Using
  Comparable Corpora (BUCC 2021)}, pages 18--27, Online (Virtual Mode). INCOMA
  Ltd.

\bibitem[{Engberg-Pedersen(1993)}]{engberg-pedersen_1993}
Elisabeth Engberg-Pedersen. 1993.
\newblock \emph{Space in Danish Sign Language: The Semantics and Morphosyntax
  of the Use of Space in a Visual Language}.
\newblock SIGNUM-Press.

\bibitem[{Forster et~al.(2014)Forster, Schmidt, Koller, Bellgardt, and
  Ney}]{forster-etal-2014-extensions}
Jens Forster, Christoph Schmidt, Oscar Koller, Martin Bellgardt, and Hermann
  Ney. 2014.
\newblock \href
  {http://www.lrec-conf.org/proceedings/lrec2014/pdf/585_Paper.pdf} {Extensions
  of the sign language recognition and translation corpus
  {RWTH}-{PHOENIX}-weather}.
\newblock In \emph{Proceedings of the Ninth International Conference on
  Language Resources and Evaluation ({LREC}'14)}, pages 1911--1916, Reykjavik,
  Iceland. European Language Resources Association (ELRA).

\bibitem[{Hanke et~al.(2020)Hanke, Schulder, Konrad, and
  Jahn}]{hanke:20016:sign-lang:lrec}
Thomas Hanke, Marc Schulder, Reiner Konrad, and Elena Jahn. 2020.
\newblock \href {https://www.sign-lang.uni-hamburg.de/lrec/pub/20016.pdf}
  {Extending the {Public} {DGS} {Corpus} in size and depth}.
\newblock In \emph{Proceedings of the {LREC2020} 9th Workshop on the
  Representation and Processing of Sign Languages: Sign Language Resources in
  the Service of the Language Community, Technological Challenges and
  Application Perspectives}, pages 75--82, Marseille, France. {European
  Language Resources Association (ELRA)}.

\bibitem[{Kocmi et~al.(2021)Kocmi, Federmann, Grundkiewicz, Junczys-Dowmunt,
  Matsushita, and Menezes}]{kocmi-etal-2021-ship}
Tom Kocmi, Christian Federmann, Roman Grundkiewicz, Marcin Junczys-Dowmunt,
  Hitokazu Matsushita, and Arul Menezes. 2021.
\newblock \href {https://aclanthology.org/2021.wmt-1.57} {To ship or not to
  ship: An extensive evaluation of automatic metrics for machine translation}.
\newblock In \emph{Proceedings of the Sixth Conference on Machine Translation},
  pages 478--494, Online. Association for Computational Linguistics.

\bibitem[{Konrad et~al.(2022)Konrad, Hanke, Langer, König, König, Nishio, and
  Regen}]{konrad_reiner_2022_10251}
Reiner Konrad, Thomas Hanke, Gabriele Langer, Susanne König, Lutz König, Rie
  Nishio, and Anja Regen. 2022.
\newblock \href {https://doi.org/10.25592/uhhfdm.10251} {{Public DGS Corpus:
  Annotation Conventions / Öffentliches DGS-Korpus: Annotationskonventionen}}.

\bibitem[{Kopf et~al.(2022)Kopf, Schulder, Hanke, and
  Bigeard}]{kopfmariaAnnotations2022}
Maria Kopf, Marc Schulder, Thomas Hanke, and Sam Bigeard. 2022.
\newblock \href {https://doi.org/10.25592/uhhfdm.9842} {Specification for the
  harmonization of sign language annotations}.

\bibitem[{Kusters et~al.(2017)Kusters, O'Brien, and {De
  Meulder}}]{kusters-et-al-2017}
{Annelies Maria Jozef} Kusters, Dai O'Brien, and Maartje {De Meulder}. 2017.
\newblock {Innovations in Deaf Studies: Critically Mapping the Field}.
\newblock In Annelies Kusters, Maartje {De Meulder}, and Dai O'Brien, editors,
  \emph{Innovations in Deaf Studies}, pages 1--53. Oxford University Press,
  United Kingdom.

\bibitem[{LeCun et~al.(2010)LeCun, Cortes, and Burges}]{lecun2010mnist}
Yann LeCun, Corinna Cortes, and CJ~Burges. 2010.
\newblock Mnist handwritten digit database.
\newblock \emph{ATT Labs [Online]. Available:
  http://yann.lecun.com/exdb/mnist}, 2.

\bibitem[{Lin(2004)}]{lin-2004-rouge}
Chin-Yew Lin. 2004.
\newblock \href {https://aclanthology.org/W04-1013} {{ROUGE}: A package for
  automatic evaluation of summaries}.
\newblock In \emph{Text Summarization Branches Out}, pages 74--81, Barcelona,
  Spain. Association for Computational Linguistics.

\bibitem[{Lipton and Steinhardt(2019)}]{lipton-steinhardt}
Zachary~C. Lipton and Jacob Steinhardt. 2019.
\newblock \href {https://doi.org/10.1145/3317287.3328534} {Troubling trends in
  machine learning scholarship: Some ml papers suffer from flaws that could
  mislead the public and stymie future research.}
\newblock \emph{Queue}, 17(1):45–77.

\bibitem[{Marie et~al.(2021)Marie, Fujita, and
  Rubino}]{marie-etal-2021-scientific}
Benjamin Marie, Atsushi Fujita, and Raphael Rubino. 2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.566} {Scientific
  credibility of machine translation research: A meta-evaluation of 769
  papers}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 7297--7306,
  Online. Association for Computational Linguistics.

\bibitem[{Morgan and Woll(2002)}]{morgan-woll-2002}
Gary Morgan and Bencie Woll. 2002.
\newblock The development of complex sentences in {British Sign Language}.
\newblock In Gary Morgan and Bencie Woll, editors, \emph{Directions in Sign
  Language Acquisition: Trends in Language Acquisition Research}, pages
  255--276. John Benjamins, Amsterdam, Netherlands.

\bibitem[{Moryossef and Goldberg(2021)}]{moryossef2021slp}
Amit Moryossef and Yoav Goldberg. 2021.
\newblock {S}ign {L}anguage {P}rocessing.
\newblock \url{https://sign-language-processing.github.io/}.

\bibitem[{Moryossef et~al.(2021)Moryossef, Yin, Neubig, and
  Goldberg}]{moryossef-etal-2021-data}
Amit Moryossef, Kayo Yin, Graham Neubig, and Yoav Goldberg. 2021.
\newblock \href {https://aclanthology.org/2021.mtsummit-at4ssl.1} {Data
  augmentation for sign language gloss translation}.
\newblock In \emph{Proceedings of the 1st International Workshop on Automatic
  Translation for Signed and Spoken Languages (AT4SSL)}, pages 1--11, Virtual.
  Association for Machine Translation in the Americas.

\bibitem[{M{\"u}ller et~al.(2018)M{\"u}ller, Rios, Voita, and
  Sennrich}]{muller-etal-2018-large}
Mathias M{\"u}ller, Annette Rios, Elena Voita, and Rico Sennrich. 2018.
\newblock \href {https://doi.org/10.18653/v1/W18-6307} {A large-scale test set
  for the evaluation of context-aware pronoun translation in neural machine
  translation}.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation:
  Research Papers}, pages 61--72, Brussels, Belgium. Association for
  Computational Linguistics.

\bibitem[{Napier and Leeson(2016)}]{napier-leeson-2016}
Jemina Napier and Lorraine Leeson. 2016.
\newblock \emph{{Sign Language in Action}}.
\newblock Palgrave Macmillan, London.

\bibitem[{Orbay and Akarun(2020)}]{orba-akarun}
Alptekin Orbay and Lale Akarun. 2020.
\newblock \href {https://doi.org/10.1109/FG47880.2020.00002} {Neural sign
  language translation by learning tokenization}.
\newblock In \emph{2020 15th IEEE International Conference on Automatic Face
  and Gesture Recognition (FG 2020)}, pages 222--228.

\bibitem[{Papineni et~al.(2002)Papineni, Roukos, Ward, and
  Zhu}]{papineni-etal-2002-bleu}
Kishore Papineni, Salim Roukos, Todd Ward, and Wei-Jing Zhu. 2002.
\newblock \href {https://doi.org/10.3115/1073083.1073135} {{B}leu: a method for
  automatic evaluation of machine translation}.
\newblock In \emph{Proceedings of the 40th Annual Meeting of the Association
  for Computational Linguistics}, pages 311--318, Philadelphia, Pennsylvania,
  USA. Association for Computational Linguistics.

\bibitem[{Pizzuto et~al.(2006)Pizzuto, Rossini, and
  Russo}]{pizzuto:06001:sign-lang:lrec}
Elena~Antinoro Pizzuto, Paolo Rossini, and Tommaso Russo. 2006.
\newblock \href {https://www.sign-lang.uni-hamburg.de/lrec/pub/06001.pdf}
  {Representing signed languages in written form: Questions that need to be
  posed}.
\newblock In \emph{Proceedings of the {LREC2006} 2nd Workshop on the
  Representation and Processing of Sign Languages: Lexicographic Matters and
  Didactic Scenarios}, pages 1--6, Genoa, Italy. {European Language Resources
  Association (ELRA)}.

\bibitem[{Popovi{\'c}(2016)}]{popovic-2016-chrf}
Maja Popovi{\'c}. 2016.
\newblock \href {https://doi.org/10.18653/v1/W16-2341} {chr{F} deconstructed:
  beta parameters and n-gram weights}.
\newblock In \emph{Proceedings of the First Conference on Machine Translation:
  Volume 2, Shared Task Papers}, pages 499--504, Berlin, Germany. Association
  for Computational Linguistics.

\bibitem[{Post(2018)}]{post-2018-call}
Matt Post. 2018.
\newblock \href {https://doi.org/10.18653/v1/W18-6319} {A call for clarity in
  reporting {BLEU} scores}.
\newblock In \emph{Proceedings of the Third Conference on Machine Translation:
  Research Papers}, pages 186--191, Brussels, Belgium. Association for
  Computational Linguistics.

\bibitem[{Saunders et~al.(2020)Saunders, Camg{\"o}z, and
  Bowden}]{saunders2020progressive}
Ben Saunders, Necati~Cihan Camg{\"o}z, and Richard Bowden. 2020.
\newblock {Progressive Transformers for End-to-End Sign Language Production}.
\newblock In \emph{Proceedings of the European Conference on Computer Vision
  (ECCV)}.

\bibitem[{Saunders et~al.(2022)Saunders, Camg{\"o}z, and
  Bowden}]{saunders2021signing}
Ben Saunders, Necati~Cihan Camg{\"o}z, and Richard Bowden. 2022.
\newblock {Signing at Scale: Learning to Co-Articulate Signs for Large-Scale
  Photo-Realistic Sign Language Production}.
\newblock In \emph{Proceedings of the IEEE Conference on Computer Vision and
  Pattern Recognition (CVPR)}.

\bibitem[{Sennrich and Zhang(2019)}]{sennrich-zhang-2019-revisiting}
Rico Sennrich and Biao Zhang. 2019.
\newblock \href {https://doi.org/10.18653/v1/P19-1021} {Revisiting low-resource
  neural machine translation: A case study}.
\newblock In \emph{Proceedings of the 57th Annual Meeting of the Association
  for Computational Linguistics}, pages 211--221, Florence, Italy. Association
  for Computational Linguistics.

\bibitem[{Stoll et~al.(2018)Stoll, Camg{\"o}z, Hadfield, and
  Bowden}]{stoll2018sign}
Stephanie Stoll, Necati~Cihan Camg{\"o}z, Simon Hadfield, and Richard Bowden.
  2018.
\newblock Sign language production using neural machine translation and
  generative adversarial networks.
\newblock In \emph{Proceedings of the 29th British Machine Vision Conference
  (BMVC 2018)}. British Machine Vision Association.

\bibitem[{Stoll et~al.(2020)Stoll, Camg{\"o}z, Hadfield, and
  Bowden}]{stoll2020text2sign}
Stephanie Stoll, Necati~Cihan Camg{\"o}z, Simon Hadfield, and Richard Bowden.
  2020.
\newblock Text2sign: towards sign language production using neural machine
  translation and generative adversarial networks.
\newblock \emph{International Journal of Computer Vision}, 128(4):891--908.

\bibitem[{Vaswani et~al.(2017)Vaswani, Shazeer, Parmar, Uszkoreit, Jones,
  Gomez, Kaiser, and Polosukhin}]{DBLP:journals/corr/VaswaniSPUJGKP17}
Ashish Vaswani, Noam Shazeer, Niki Parmar, Jakob Uszkoreit, Llion Jones,
  Aidan~N Gomez, {\L}ukasz Kaiser, and Illia Polosukhin. 2017.
\newblock {Attention is All you Need}.
\newblock In \emph{{Advances in Neural Information Processing Systems 30}},
  pages 5998--6008.

\bibitem[{Walsh et~al.(2022)Walsh, Saunders, and
  Bowden}]{walsh-saunders-bowden:2022:SLTAT}
Harry Walsh, Ben Saunders, and Richard Bowden. 2022.
\newblock \href {https://aclanthology.org/2022.sltat-1.18} {Changing the
  representation: Examining language representation for neural sign language
  production}.
\newblock In \emph{Proceedings of the 7th International Workshop on Sign
  Language Translation and Avatar Technology: The Junction of the Visual and
  the Textual: Challenges and Perspectives}, pages 117--124, Marseille, France.
  European Language Resources Association.

\bibitem[{Yin et~al.(2021)Yin, Moryossef, Hochgesang, Goldberg, and
  Alikhani}]{yin-etal-2021-including}
Kayo Yin, Amit Moryossef, Julie Hochgesang, Yoav Goldberg, and Malihe Alikhani.
  2021.
\newblock \href {https://doi.org/10.18653/v1/2021.acl-long.570} {Including
  signed languages in natural language processing}.
\newblock In \emph{Proceedings of the 59th Annual Meeting of the Association
  for Computational Linguistics and the 11th International Joint Conference on
  Natural Language Processing (Volume 1: Long Papers)}, pages 7347--7360,
  Online. Association for Computational Linguistics.

\bibitem[{Yin and Read(2020)}]{yin-read-2020-better}
Kayo Yin and Jesse Read. 2020.
\newblock \href {https://doi.org/10.18653/v1/2020.coling-main.525} {Better sign
  language translation with {STMC}-transformer}.
\newblock In \emph{Proceedings of the 28th International Conference on
  Computational Linguistics}, pages 5975--5989, Barcelona, Spain (Online).
  International Committee on Computational Linguistics.

\bibitem[{Zhang and Duh(2021)}]{zhang-duh-2021-approaching}
Xuan Zhang and Kevin Duh. 2021.
\newblock \href {https://aclanthology.org/2021.mtsummit-at4ssl.7} {Approaching
  sign language gloss translation as a low-resource machine translation task}.
\newblock In \emph{Proceedings of the 1st International Workshop on Automatic
  Translation for Signed and Spoken Languages (AT4SSL)}, pages 60--70, Virtual.
  Association for Machine Translation in the Americas.

\end{thebibliography}

\pdfoutput=1

\documentclass[11pt]{article}

\usepackage{EMNLP2022}

\usepackage{times}
\usepackage{latexsym}

\usepackage[T1]{fontenc}

\usepackage[utf8]{inputenc}

\usepackage{microtype}

\usepackage{inconsolata}

\usepackage{multirow}
\usepackage{booktabs}
\usepackage{pifont}
\usepackage[frozencache,cachedir=minted-cache]{minted}

\title{Considerations for meaningful sign language machine translation based on glosses}

\author{Mathias M\"{u}ller$^1$, Zifan Jiang$^1$, Amit Moryossef$^{1,2,3}$, Annette Rios$^1$ and Sarah Ebling$^1$ \\
  $^1$ Department of Computational Linguistics, University of Zurich, Switzerland \\
  $^2$ ETH Zurich, Switzerland, $^3$ Bar-Ilan University, Tel Aviv, Israel \\
  \texttt{\{mmueller,jiang,rios,ebling\}@cl.uzh.ch}, \texttt{amitmoryossef@gmail.com}}

\begin{document}
\maketitle
\begin{abstract}

Automatic sign language processing is gaining popularity in Natural Language Processing (NLP) research \citep{yin-etal-2021-including}. In machine translation (MT) in particular, sign language translation based on \textit{glosses} is a prominent approach.
In this paper, we review recent works on neural gloss translation. We find that limitations of glosses in general and limitations of specific datasets are not discussed in a transparent manner and that there is no common standard for evaluation.

To address these issues, we put forward concrete recommendations for future research on gloss translation. Our suggestions advocate awareness of the inherent limitations of gloss-based approaches, realistic datasets, stronger baselines and convincing evaluation.

\end{abstract}

\section{Introduction}
\label{sec:intro}

Automatic sign language processing is becoming more popular in Natural Language Processing (NLP) research \citep{yin-etal-2021-including}. In machine translation (MT) in particular, many recent publications have proposed sign language translation based on \textit{glosses}. Glosses provide semantic labels for individual signs. They typically consist of the base form of a word in the surrounding spoken language written in capital letters (see Table \ref{tbl:gloss_example}).
Even though glosses are not a complete representation of signs (see e.g. \citealt{pizzuto:06001:sign-lang:lrec}), they are often adopted in MT because, by virtue of being textual, they fit seamlessly into existing MT pipelines and existing methods seemingly require the least modification.

In this paper, we review recent works on neural gloss translation. We find that limitations of gloss-based approaches in general and limitations of specific datasets are not transparently discussed as inherent shortcomings. Furthermore, among gloss translation papers there is no common standard for evaluation, especially regarding the exact method to compute BLEU scores.


Experiments in sign language translation should be informed by sign language expertise and should be performed according to the best practices already established in the MT community.

To alleviate these problems going forward, we make practical recommendations for future research on gloss translation.

\begin{table}
    \centering
    \small
    \begin{tabular}{p{0.9\linewidth}}
      \toprule
      \textbf{Glosses (DSGS)} \\
      \texttt{KINDER FREUEN WARUM FERIEN NÄHER-KOMMEN} \\
      \midrule
      \textbf{Translation (DE)} \\
      \texttt{Die Kinder freuen sich, weil die Ferien näher rücken.} \\
      \midrule
      \textbf{Glosses (EN)} \\
      \texttt{(`CHILDREN REJOICE WHY HOLIDAYS APPROACHING')} \\
      \midrule
      \textbf{Translation (EN)} \\
      \texttt{(`The children are happy because the holidays are approaching.')} \\
      \bottomrule
    \end{tabular}
    \caption{Example for sign language glosses. DSGS=Swiss German Sign Language, DE=German, EN=English. English translations are provided for convenience. Example is adapted from a lexicon of the three sign languages of Switzerland, where a sign language video of this sentence is available (\url{https://signsuisse.sgb-fss.ch/de/lexikon/g/ferien/}).}
    \label{tbl:gloss_example}
\end{table}



Our paper makes the following contributions:

\begin{itemize}
    \item We provide a review of recent works on gloss translation (\S\ref{sec:related_work}).
    \item We outline recommendations for future work which promote awareness of the inherent limitations of gloss-based approaches, realistic datasets, stronger baselines and convincing evaluation (\S\ref{sec:best_practices_for_gloss_translation}).
\end{itemize}

\section{Related work}
\label{sec:related_work}


\begin{table*}
    \centering
    \tiny
    \begin{tabular}{l c cc ccc c cccc l}
      \toprule
      \multicolumn{1}{c}{} & 
      \multicolumn{1}{c}{\textbf{L}} & 
      \multicolumn{2}{c}{\textbf{datasets}} & \multicolumn{3}{c}{\textbf{translation directions}} & \multicolumn{1}{c}{\textbf{code}} & \multicolumn{4}{c}{\textbf{evaluation metrics}} & 
      \multicolumn{1}{c}{\textbf{BLEU tool}} \\
      \cmidrule(lr){2-2}
      \cmidrule(lr){3-4}
      \cmidrule(lr){5-7}
      \cmidrule(lr){8-8}
      \cmidrule(lr){9-12}
      \cmidrule(lr){13-13}
      & & \textbf{P} & \textbf{O} & \textbf{DGS$\rightarrow$DE} & \textbf{DE$\rightarrow$DGS} & \textbf{O} & & \textbf{B 1-3} & \textbf{B-4} & \textbf{R} & \textbf{O} \\
      
      \midrule
      
      
      \citet{8578910} & - & \ding{52} & - & \ding{52} & - & - & \href{https://github.com/neccam/nslt}{\ding{52}} & \ding{52} & \ding{52} & \ding{52} & - & \href{https://github.com/tensorflow/nmt/blob/master/nmt/scripts/bleu.py}{Tensorflow}
       \\
       
       \midrule
      
      
      \citet{stoll2018sign} & - & \ding{52} & - & - & \ding{52} & - & - & \ding{52} & \ding{52} & \ding{52} & - & (unclear) \\
      
      \midrule
      
      \citet{camgoz2020sign} & \ding{52} & \ding{52} & - & \ding{52} & - & - & \href{https://github.com/neccam/slt}{\ding{52}} & \ding{52} & \ding{52} & - & WER & (unclear)
       \\
      
      \midrule
      
      \citet{camgoz2020b} & \ding{52} & \ding{52} & - & \ding{52} & - & - & - & - & \ding{52} & \ding{52} & - & (unclear)
       \\
      
      \midrule
      
      \citet{yin-read-2020-better} & \ding{52} & \ding{52} & ASLG-PC12 & \ding{52} & - & ASL$\rightarrow$EN & \href{https://github.com/kayoyin/transformer-slt}{\ding{52}} & \ding{52} & \ding{52} & \ding{52} & METEOR & \href{https://github.com/kayoyin/transformer-slt/blob/master/tools/bleu.py}{NLTK} \\
      
      \midrule
      
      \citet{saunders2020progressive} & \ding{52} & \ding{52} & - & - & \ding{52} & - & \href{https://github.com/BenSaunders27/ProgressiveTransformersSLP}{\ding{52}} & \ding{52} & \ding{52} & \ding{52} & - & (unclear)
       \\
      
      \midrule
      
      \citet{stoll2020text2sign} & \ding{52} & \ding{52} & - & - & \ding{52} & - & - & \ding{52} & \ding{52} & \ding{52} & WER & (unclear)  \\
      
      \midrule
      
      \citet{orba-akarun} & - & \ding{52} & - & \ding{52} & - & - & - & \ding{52} & \ding{52} & \ding{52} & - & (unclear) \\
      
      \midrule
      
      \citet{moryossef-etal-2021-data} & - & \ding{52} & NCSLGR & \ding{52} & - & ASL$\rightarrow$EN & (\ding{52}) & - & \ding{52} & - & COMET & \href{https://github.com/mjpost/sacrebleu}{SacreBLEU} \\
      
      \midrule
      
      \citet{zhang-duh-2021-approaching} & - & \ding{52} & - & \ding{52} & \ding{52} & - & - & - & \ding{52} & - & - & (unclear) \\
      
      \midrule
      
      \citet{egea-gomez-etal-2021-syntax} & - & \ding{52} & - & - & \ding{52} & - & \href{https://github.com/LaSTUS-TALN-UPF/Syntax-Aware-Transformer-Text2Gloss}{\ding{52}} & - & \ding{52} & \ding{52} & METEOR, TER & \href{https://github.com/LaSTUS-TALN-UPF/Syntax-Aware-Transformer-Text2Gloss/blob/main/utils/sacreBLEU_script.py}{SacreBLEU}  \\
      
      \midrule
      
      \citet{saunders2021signing} & - & \ding{52} & DGS Corpus & - & \ding{52} & - & - & - & \ding{52} & \ding{52} & - & (unclear) \\
      
      \midrule
      
      \citet{angelova-etal-2022-using} & \ding{52} & \ding{52} & DGS Corpus & \ding{52} & - & - & \href{https://github.com/DFKI-SignLanguage/gloss-to-text-sign-language-translation}{\ding{52}} & - & \ding{52} & - & - & SacreBLEU \\
      
      \midrule
      
      \citet{walsh-saunders-bowden:2022:SLTAT} & - & \ding{52} & DGS Corpus & - & \ding{52} & - & - & \ding{52} & \ding{52} & \ding{52} & - &  (unclear) \\
      
      \bottomrule
    \end{tabular}
    \caption{Review of recent works on gloss translation. L=whether a paper discusses limitations of gloss approaches, P=\texttt{RWTH-PHOENIX-Weather 2014T} introduced by \citet{8578910}, O=other, (\ding{52})=exact shell commands are listed in the appendix of the paper, B=BLEU, R=ROUGE, B1-4=four variants of BLEU, varying the maximum ngram order from 1 to 4, SacreBLEU=version 1.4.14, COMET=\href{https://unbabel.github.io/COMET/html/models.html}{wmt-large-da-estimator-1719}, DGS=German Sign Language, DE=German, ASL=American Sign Language, EN=English. In the code column, checkmark symbols (\ding{52}) are clickable links.}
    \label{tbl:literature_review}
\end{table*}




For a general overview of sign language processing in the context of NLP see \citet{yin-etal-2021-including,moryossef2021slp} and \citet{de2022machine} for a comprehensive survey of sign language machine translation (including, but not limited to, gloss-based approaches).

We conduct a more narrow literature review of 14 recent publications on gloss translation. We report characteristics such as the datasets used, translation directions, and evaluation details (Table \ref{tbl:literature_review}).


\subsection{Awareness of limitations of gloss approach}
\label{subsec:awareness_of_limitations_of_gloss_approach}


We find that 8 out of 14 reviewed works do not include an adequate discussion of the limitations of gloss approaches, inadvertently overstating the potential usefulness of their experiments.

In the context of sign languages,
glosses are unique identifiers for individual signs. However, a linear sequence of glosses is not an adequate representation of a signed utterance, where different channels (manual and non-manual) are engaged simultaneously. Linguistically relevant cues such as non-manual movement or use of three-dimensional space may be missing \citep{yin-etal-2021-including}.

The gloss transcription conventions of different corpora vary greatly, as does the level of detail (see \citet{kopfmariaAnnotations2022} for an overview of differences and commonalities between corpora).
Therefore, glosses in different corpora or across languages are not comparable. Gloss transcription is an enormously laborious process done by expert linguists.

Besides, glosses are a linguistic tool, not a writing system established in Deaf
\footnote{It is a widely recognized convention to use the uppercased word \emph{Deaf} for describing members of a sign language community and, in contrast, to use the lowercased word \emph{deaf} when describing the audiological state of a hearing loss \citep{morgan-woll-2002}. More recent works \citep{napier-leeson-2016,kusters-et-al-2017} propose to use \emph{deaf} in both situations.}
communities. Sign language users generally do not read or write glosses in their everyday lives.

Taken together, this means that gloss translation suffers from an inherent and irrecoverable information loss,
that creating an abundance of translations transcribed as glosses is unrealistic, and that gloss translation systems are not immediately useful to end users.

\subsection{Choice of dataset}
\label{subsec:choice_of_dataset}

All reviewed works use the \texttt{RWTH-PHOENIX Weather 2014T} (hereafter abbreviated as \textit{PHOENIX}) dataset \citep{forster-etal-2014-extensions,8578910} while other datasets are used far less frequently. Besides, we note a distinct paucity of languages and translation directions: 12 out of 14 works are concerned only with translation between German Sign Language (DGS) and German (DE), the language pair of the PHOENIX dataset.

While PHOENIX was a breakthrough when it was published, it is of limited use for current research. 
The dataset is small (8k sentence pairs) and contains only weather reports, covering a very narrow linguistic domain. It is important to discuss the exact nature of glosses, how the corpus was created and how it is distributed.

\paragraph{Glossing}
PHOENIX is based on German weather reports interpreted into DGS and broadcast on the TV station Phoenix.
The broadcast videos served as input for the DGS side of the parallel corpus. Compared to the glossing conventions of other well-known corpora, PHOENIX glosses are simplistic and capture mostly manual features (with mouthings as the only non-manual activity covered), which is not sufficient to represent meaning (\S \ref{subsec:awareness_of_limitations_of_gloss_approach}).

\paragraph{Live interpretation and translationese effects} The fact that PHOENIX data comes from interpretation in a live setting has two implications: Firstly, since information was conveyed at high speed, the sign language interpreters omitted pieces of information from time to time. This leads to an information mismatch between some German sentences and their DGS counterparts. Secondly, due to the high speed of transmission, the (hearing) interpreters sometimes followed the grammar of German more closely than that of DGS. As a consequence, some translations are more similar to Signed German (\textit{Lautsprachbegleitendes Gebärden}) than DGS. 

\paragraph{Preprocessing of spoken language} The German side of the PHOENIX corpus is available only already tokenized, lowercased and with punctuation symbols removed. From an MT perspective this is unexpected since corpora are usually distributed without such preprocessing.

PHOENIX is popular because it is freely available and is a benchmark with clearly defined data splits introduced by \citet{8578910}. Sign language machine translation as a field is experiencing a shortage of free and open datasets and, with the exception of PHOENIX, there are no agreed-upon data splits.


Essentially, from a scientific point of view achieving higher gloss translation quality on the PHOENIX dataset is near meaningless. The apparent overuse of PHOENIX is reminiscent of the overuse of MNIST \citep{lecun2010mnist} in machine learning, or the overuse of the WMT 14 English-German benchmark in the MT community, popularized by \citet{DBLP:journals/corr/VaswaniSPUJGKP17}.

\subsection{Evaluation}
\label{subsec:evaluation}

As evaluation metrics, all works use some variant of BLEU \citep{papineni-etal-2002-bleu}, and ten out of 14 use some variant of ROUGE \citep{lin-2004-rouge}. All but four papers do not contain enough information about how exactly BLEU was computed. Different BLEU implementations, settings (e.g. ngram orders, tokenization schemes) and versions are used.


\begin{table*}
    \centering
    \small
    \begin{tabular}{lp{10cm}}
      \toprule
      \textbf{Reference} & \texttt{VIEL1A FAMILIE1* JUNG1 FAMILIE1 GERN1* IN1* HAMBURG1* STADT2* WOHNUNG2B* FAMILIE1} \\
      \textbf{Hypothesis} & \texttt{VIEL1B JUNG1 LEBEN1 GERN1* HAMBURG1* STADT2* \$INDEX1} \\
      \midrule
      \textbf{BLEU with tokenization} & 25.61 \\
      \textbf{BLEU without tokenization} & 10.18 \\
      \bottomrule
    \end{tabular}
    \caption{Impact of applying or disabling internal tokenization (mtv13a) when computing BLEU on gloss outputs. Example taken from the Public DGS Corpus \citep{hanke:20016:sign-lang:lrec}.}
    \label{tbl:bleu_tokenization_difference}
\end{table*}

\paragraph{Non-standard metrics} ROUGE is a metric common in automatic summarization but not in MT, and was never correlated with human judgement in a large study.
In eight out of 14 papers, BLEU is used with a non-standard maximum ngram order, producing variants such as BLEU-1, BLEU-2, etc. Similar to ROUGE, these variants of BLEU have never been validated as metrics of translation quality, and their use is scientifically unmotivated.


\paragraph{Tokenization}

BLEU requires tokenized machine translations and references. Modern tools therefore apply a tokenization procedure internally and implicitly (independently of the MT system's preprocessing).
Computing BLEU with tokenization on glosses leads to seemingly better scores but is misleading since tokenization creates many trivial matches.
For instance, in corpora that make use of the character \texttt{\$} in glosses (e.g. the DGS Corpus \citep{konrad_reiner_2022_10251}), \texttt{\$} is split off as a single character, inflating the ngram sub-scores.
For an illustration see Table \ref{tbl:bleu_tokenization_difference} (and Appendix \ref{appendix:impact_of_internal_tokenization} for a complete code listing) where we demonstrate that using or omitting tokenization leads to a difference of 15 BLEU.

\paragraph{Spurious gains}

Different implementations of BLEU or different tokenizations lead to differences in BLEU bigger than what many papers describe as an ``improvement'' over previous work \citep{post-2018-call}. Incorrectly attributing such improvements to, for instance, changes to the model architecture amounts to a ``failure to identify the sources of empirical gains'' \citep{lipton-steinhardt}.
In a similar vein, we observe that papers on gloss translation tend to copy scores from previous papers without knowing whether the evaluation procedures are in fact the same. This constitutes a general trend in recent MT literature \citep{marie-etal-2021-scientific}.

In summary, some previous works on gloss translation have used 1) automatic metrics that are not suitable for MT or 2)
well-established MT metrics in ways that are not recommended.
BLEU with standard settings and tools is inappropriate for gloss outputs.

The recommended way to compute BLEU on gloss output is to use the tool SacreBLEU \citep{post-2018-call} and to disable internal tokenization.
Nevertheless, even with these precautions, it is important to note that BLEU was never validated empirically as an evaluation metric for gloss output. Some aspects of BLEU may not be adequate for a sequence of glosses, such as its emphasis on whitespaces to mark the boundaries of meaningful units that are the basis of the final score.

Other string-based metrics such as CHRF \citep{popovic-2016-chrf} may be viable alternatives for gloss evaluation. CHRF is a character-based metric and its correlation with human judgement is at least as good as BLEU's \citep{kocmi-etal-2021-ship}.

\subsection{Further observations}
\label{subsec:further_observations}

More informally (beyond what we show in Table \ref{tbl:literature_review}), we observe that most papers do not process glosses in any corpus-specific way, that particular modeling and training decisions may not be ideal for low-resource gloss translation and that document-level systems may be crucial.


\paragraph{Preprocessing glosses}

Glosses are created for linguistic purposes (\S \ref{subsec:awareness_of_limitations_of_gloss_approach}), not necessarily with machine translation in mind. Particular gloss parts are not relevant for translation and, if kept, make the problem harder unnecessarily.
For instance, a corpus transcription and annotation scheme might prescribe that meaning-equivalent, minor form  variants of signs\footnote{In spoken language linguistics, these occurrences are called \emph{allophonic variants}. We refrain from using the term here, as sign languages do not consist of sounds (\emph{phones}).} are transcribed as different glosses.
An example of meaning-equivalent form variants in some sign languages is a flat handshape with the thumb adducted (aligned with the remaining fingers) or spread.
%
Reading or producing such meaning-equivalent variants may not be relevant for an MT system and makes the learning problem harder.

Since the particular nature of glosses is specific to every corpus, it is necessary to preprocess glosses in a corpus-specific way. We illustrate corpus-specific gloss processing in Appendix \ref{appendix:example_for_corpus_specific_gloss_preprocessing}, using the Public DGS Corpus \cite{hanke:20016:sign-lang:lrec} as an example.
 

\paragraph{Modeling and training decisions}


Gloss translation experiments are certainly low-resource scenarios and therefore, best practices for optimizing MT systems on low-resource datasets apply \citep{sennrich-zhang-2019-revisiting}. For example, dropout rates or label smoothing should be set accordingly, and the vocabulary of subwords of a subword model should be generally small \citep{ding-etal-2019-call}.


Gloss translation models are often compared to other approaches as baselines, it is therefore problematic if those gloss baselines are weak and unoptimized \citep{denkowski-neubig-2017-stronger}.

\paragraph{Limitations of sentence-level systems}

All surveyed works propose sentence-level systems, as opposed to larger-context models.
For sign language MT specifically, sentence-level approaches have inherent limitations.

Signed languages exhibit frequent referencing (or ``indexing'') behaviour similar to anaphora in spoken languages. But while in spoken languages ambiguities arising from anaphora can often be resolved with sentence-level context, that is not the case for signed languages. After a subject was introduced with a sign once it is often unacceptable to use the sign again in a later utterance of the discourse, instead the subject is referred to with an index (see e.g. \citealp{engberg-pedersen_1993}). Besides, sign languages are generally not gendered (which for spoken languages helps disambiguate by narrowing down
possible antecedents).


Document-level benchmarks (e.g. \citealp{bawden-etal-2018-evaluating,muller-etal-2018-large}) show that for spoken languages, document-level context is not crucial since sentence-level baseline systems have relatively high performance. For signed languages, discourse context may be indispensable.

We are not aware of any existing discourse-level translation system for sign languages but we believe they could be investigated in the future.

\section{Recommendations for gloss translation}
\label{sec:best_practices_for_gloss_translation}

Based on our review of recent works on gloss translation, we make the following recommendations for future research:

\begin{itemize}
    \item Demonstrate awareness of limitations of gloss approaches (\S\ref{subsec:awareness_of_limitations_of_gloss_approach}), and explicitly discuss them in the paper.
    \item Focus on datasets beyond PHOENIX. Openly discuss the limited size and linguistic domain of PHOENIX (\S\ref{subsec:choice_of_dataset}).
    \item Use metrics that are well-established in MT. Compute BLEU with SacreBLEU, report metric signatures and disable internal tokenization for gloss outputs. Do not compare to scores produced with a different or unknown evaluation procedure (\S\ref{subsec:evaluation}).
    \item Given that glossing is corpus-specific (\S\ref{subsec:awareness_of_limitations_of_gloss_approach}), process glosses in a corpus-specific way, informed by the respective transcription conventions (\S\ref{subsec:further_observations}).
    \item Optimize gloss translation baselines with methods shown to be effective for low-resource MT (\S\ref{subsec:further_observations}).
\end{itemize}

We also believe publishing reproducible code makes works on gloss translation more valuable. 

\section{Conclusion}

In this paper we have shown that some recent works on gloss translation lack awareness of the inherent limitations of glosses and common datasets, as well as a standardized evaluation method (\S\ref{sec:related_work}). In order to make future research on gloss translation more meaningful, we make practical recommendations (\S\ref{sec:best_practices_for_gloss_translation}).

We urge researchers to spell out limitations of gloss translation approaches, e.g. in the now mandatory limitation sections of *ACL papers, and to strengthen their findings by implementing existing best practices in MT.

Finally, we also caution that researchers should consider whether gloss translation is worthwhile, and if time and effort would be better spent on basic linguistic tools (such as segmentation, alignment or coreference resolution), creating training corpora or translation methods that do not rely on glosses.











\section*{Data licensing}

The license of the Public DGS Corpus\footnote{\url{https://www.sign-lang.uni-hamburg.de/meinedgs/ling/license_en.html}} (which we use only as examples in Table \ref{tbl:bleu_tokenization_difference} and Appendix \ref{appendix:example_for_corpus_specific_gloss_preprocessing}) does not allow any computational research except if express permission is given by the University of Hamburg.

\section*{Acknowledgements}

This work was funded by the EU Horizon 2020 project EASIER (grant agreement no. 101016982) and the Swiss Innovation Agency  (Innosuisse) flagship IICT (PFFS-21-47).

We thank the DGS Corpus team at the University of Hamburg for helpful discussions on gloss preprocessing.

\bibliography{custom}
\bibliographystyle{acl_natbib}

\clearpage
\onecolumn

\appendix

\section{Impact of internal tokenization when computing BLEU on gloss sequences}
\label{appendix:impact_of_internal_tokenization}

\begin{listing}[!ht]
\begin{minted}[linenos, frame=lines, xleftmargin=20pt, fontsize=\small]{python}

# ! pip install sacrebleu==2.2.0

>>> from sacrebleu.metrics import BLEU

# English translation: Many young families like living in the city of Hamburg.
# German translation: Viele junge Familien leben gerne in Hamburg in der Stadt.

>>> ref = "VIEL1A FAMILIE1* JUNG1 FAMILIE1 GERN1* IN1* HAMBURG1* STADT2* WOHNUNG2B* FAMILIE1"

>>> hyp = "VIEL1B JUNG1 LEBEN1 GERN1* HAMBURG1* STADT2* $INDEX1"

# computing BLEU on gloss output with tokenization (not recommended):

>>> bleu = BLEU() # default: BLEU(tokenize="13a")
>>> bleu.corpus_score([hyp], [[ref]])
BLEU = 25.61 63.6/50.0/33.3/25.0 (BP = 0.635 ratio = 0.688 hyp_len = 11 ref_len = 16)

# computing BLEU on gloss output without tokenization (recommended):

>>> bleu = BLEU(tokenize="none")
>>> bleu.corpus_score([hyp], [[ref]])
BLEU = 10.18 57.1/16.7/10.0/6.2 (BP = 0.651 ratio = 0.700 hyp_len = 7 ref_len = 10)

\end{minted}
\caption{Impact of enabling or disabling internal tokenization (mtv13a) when computing BLEU on gloss outputs.}
\label{listing:impact_of_incorrectly_applying_internal_tokenization}
\end{listing}

\section{Example for corpus-specific gloss preprocessing}
\label{appendix:example_for_corpus_specific_gloss_preprocessing}

For this example we recommend to download and process release 3.0 of the corpus.
To DGS glosses we suggest to apply the following modifications derived from the DGS Corpus transcription conventions \citep{konrad_reiner_2022_10251}:

\begin{itemize}
    \item Removing entirely two specific gloss types that cannot possibly help the translation: \texttt{\$GEST-OFF} and \texttt{\$\$EXTRA-LING-MAN}.
    \item Removing \textit{ad-hoc} deviations from citation forms, marked by \texttt{*}. Example: \texttt{ANDERS1*} $\rightarrow$ \texttt{ANDERS1}.
    \item Removing the distinction between type glosses and subtype glosses, marked by \texttt{\^{}}. Example: \texttt{WISSEN2B\^{}} $\rightarrow$ \texttt{WISSEN2B}. 
    \item Collapsing phonological variations of the same type that are meaning-equivalent. Such variants are marked with uppercase letter suffixes. Example: \texttt{WISSEN2B} $\rightarrow$ \texttt{WISSEN2}.
    \item Deliberately keep numerals (\texttt{\$NUM}), list glosses (\texttt{\$LIST}) and finger alphabet (\texttt{\$ALPHA}) intact, except for removing handshape variants.
\end{itemize}

See Table \ref{tbl:gloss_preprocessing_examples} for examples for this preprocessing step. Overall these simplifications should reduce the number of observed forms while not affecting the machine translation task. For other purposes such as linguistic analysis our preprocessing would of course be detrimental.

\newpage

\begin{table*}
    \centering
    \small
    \begin{tabular}{lp{13cm}}
      \toprule
      \textbf{before} & \texttt{\$INDEX1 ENDE1\^{} ANDERS1* SEHEN1 MÜNCHEN1B* BEREICH1A*} \\
      \textbf{after} & \texttt{\$INDEX1 ENDE1 ANDERS1 SEHEN1 MÜNCHEN1 BEREICH1} \\
      \midrule
      \textbf{before} & \texttt{ICH1 ETWAS-PLANEN-UND-UMSETZEN1 SELBST1A* KLAPPT1* \$GEST-OFF\^{} BIS-JETZT1 GEWOHNHEIT1* \$GEST-OFF\^{}*} \\
      \textbf{after} & \texttt{ICH1 ETWAS-PLANEN-UND-UMSETZEN1 SELBST1 KLAPPT1 BIS-JETZT1 GEWOHNHEIT1} \\
      \bottomrule
    \end{tabular}
    \caption{Examples for preprocessing of DGS glosses.}
    \label{tbl:gloss_preprocessing_examples}
\end{table*}

While this preprocessing method provides a good baseline, it can certainly be refined further. For instance, the treatment of two-handed signs could be improved. If a gloss occurs simultaneously on both hands, we either keep both glosses or remove one occurrence. In both cases, information about the simultaneity of signs is lost during preprocessing and preserving it could potentially improve translation.

\end{document}